Air streams horizontally past a small airplane's w
Answer #1

Here ,

v1 = 70 m/s

v2 = 60 m/s

density of air , p = 1.20 Kg/m^3

area . A = 16.2 m^2

Pressure difference = 0.5 * p * (v1^2 - v2^2)

Pressure difference = 0.5 * 1.20 * (70^2 - 60^2)

Pressure difference = 780 Pa

the pressure difference on the wing is 780 Pa

--------

net lift force = pressure * area

net lift force = 780 * 16.2

net lift force = 12636 N

No , other data is required